# Searching for Free-Floating Planets with TESS: I. Discovery of a First Terrestrial-Mass Candidate

### v1

@@ 3.3 Automated Flux-Level Vetting
A total of 58,303 signals met our basic detection criteria and were sent through a first round of automated vetting, which involved computing a set of metrics to be compared to pass-fail thresholds. If a signal failed to meet any of our criteria, it was rejected as a possible microlensing candidate. We do not claim that these empirically-determined thresholds are optimal and, as is the case for any vetting procedure, they may have removed true signals as well as false positives. However, they successfully produced a dataset of microlensing candidates amenable to manual inspection (§3.5).
@@ 3.3.1 Initial Model Fitting
Each signal was first fit to three different models: a straight line parameterized by a slope and offset; a skew normal distribution parameterized by a central timescale, amplitude, scale ($\omega$), and shape ($\alpha$); and a simple point-source lens model parameterized by $t_{0}$, $u_{0}$, and $t_{E}$. Fitting was performed using the lmfit Python package (Newville et al., 2016). Data more than $2\tau$ from $t_{\mathrm{BLS}}$ were ignored in order to speed up the fitting process.
A candidate was rejected if a point lens was a worse fit compared to a straight line according to the reduced $\chi^{2}$ of each fit, i.e., $\chi^{2}_{\mathrm{lens}}>\chi^{2}_{\mathrm{line}}$. Only 526 signals failed this criteria; however, the fit results were more useful for other tests targeted at specific types of false positives.
@@ 3.3.2 Edge Effects and Gapped Events
BLS occasionally triggers at the edge of data gaps, especially if such edges correspond to rapid increases in brightness due to scattered light, a well-known systematic effect within TESS data. A number of events also lack a significant amount of data either during the brightening, or just before and/or after. To identify these cases, we computed the fraction of the light curve near the event that actually contained data. We found $n_{\mathrm{exp}}$, the number of data points expected within $2\tau$ of $t_{\mathrm{BLS}}$ given the 200-second cadence of the observations, as well as $n_{\mathrm{obs}}$, the number of data points actually observed over that time. We rejected 46,117 signals with $n_{\mathrm{obs}}/n_{\mathrm{exp}}<0.9$. Given that this cut rejected 79% of all vetted signals, edge effects and gapped events constituted the majority of TESS microlensing false positives.
@@ 3.3.3 Non-Unique Events
Microlensing events are rare, and should therefore be unique compared to other events in the light curve with the same duration. The presence of similar features at other times in a light curve suggest short-term periodic or quasi-periodic stellar variability, or low-SNR noise/instrumental systematics. To quantify uniqueness, we adopted elements from the Model-Shift Uniqueness Test developed for Kepler transiting exoplanet candidate vetting (Coughlin, 2017). In summary, we computed an SNR time series by finding the SNR of events centered at each point in the light curve over the duration of the BLS-detected signal. We measured the significance of the primary microlensing event ($\mathrm{SNR}_{\mathrm{pri}}$) as the largest SNR value in the SNR time series within a half event duration of the BLS detection. The most significant event at least two event durations from the primary was labeled the secondary event ($\mathrm{SNR}_{\mathrm{sec}}$), and the next most significant event at least two event durations from both the primary and secondary was labeled the tertiary event ($\mathrm{SNR}_{\mathrm{ter}}$). Finally, the most significant negative event ($\mathrm{SNR}_{\mathrm{neg}}$), representing the largest flux decrease at least three event durations from the primary and secondary, was also labeled.
We rejected signals for which the primary event was not sufficiently unique, as quantified by $\mathrm{SNR}_{\mathrm{pri}}<10$ or $\Delta\mathrm{SNR}<2$ compared to all other events (the secondary, tertiary, or most negative events). A total of 6,797 signals failed this criteria.
@@ 3.3.4 Flares
Flares are highly asymmetric brightening events characterized by a rapid increase in flux followed by a gradual decay, in contrast to the symmetric, bell-shaped peaks of microlensing events. We used the results of the skew normal distribution fit to identify this class of false positives. We rejected 41,615 signals with a skew shape of $|\alpha|>2$, suggesting strong asymmetry. Half of these also failed the edge effect/gapped event test, indicating that they were likely asymmetric events caused by scattered light. We also rejected 2,152 signals with a skew scale of $\omega<0.01$ days, as these are consistent with extremely short-duration flares or single-cadence events such as cosmic ray hits.
@@ 3.4 Asteroid Rejection
A total of 6,524 signals passed all previous tests, and the most common remaining false positive scenario was brightening events caused by passing asteroids. We employed three major tests to reject asteroids: (1) cross-matching our candidates with known objects in the Solar System, (2) analyzing each event at the pixel level, and (3) searching for indications of moving bodies by matching correlated events across RA and Dec space.
@@
(1) Known asteroids: We queried the NASA/JPL Small-Body Identification API [^2] https://ssd-api.jpl.nasa.gov/doc/sb_ident.html for all known small bodies seen by TESS within the entire region in RA and Dec covered by Sector 61 in each camera and CCD combination, extracting all objects brighter than 21 mag seen between 2962.79 and 2990.29 TESS BJD in steps of 2.5 days. We found 8,270 unique objects. Ephemerides for each were generated in 30-minute intervals using the JPL Horizons Solar System data and ephemeris computation service [^3] https://ssd.jpl.nasa.gov/horizons/. We then calculated the angular separation and difference in time between each asteroid and each surviving signal. We rejected 6,124 signals with angular separation of $<0.07^{\circ}$ and time difference of $<0.2$ days.
@@
(2) Pixel-level analysis: Hundreds of asteroids were not rejected by this automated cut due to incompleteness of the Small-Body Database. We further removed asteroids by inspecting each candidate at the pixel level. Using $21\times 21$ pixel cutouts centered on each target star produced by tesscut (Brasseur et al., 2019), we computed the mean of all out-of-event images to represent the typical field near a source star when no brightening or dimming events are occurring on average. Out-of-event times were defined as those between $0.5\tau$ and $1.5\tau$ from the center of each event. We then subtracted the mean out-of-event image from each individual image in order to find the pixels undergoing the largest brightening at each point in time. We searched for changes in the locations of the brightest pixels within $1.5\tau$ of the event, which are consistent with an asteroid moving across the TESS pixels. Examples showing the maximum values of each pixel, for both an asteroid and a likely isolated candidate, are provided in Figure 1, highlighting the clear difference between the two scenarios.
@@
Figure 1: Pixel-level data demonstrating typical signatures for asteroids (left) compared to isolated events (right), using examples associated with TIC-141479395 and TIC-432012378, respectively. These data were produced by subtracting the mean out-of-event image from each image within $1.5\tau$ of the center of a given candidate event, and finding the maximum value of each pixel.
@@
We also produced difference images, employing the same techniques commonly used for transiting exoplanet vetting (Bryson et al., 2013), with the TESS difference imaging codebase transit-diffImage. [^4] https://github.com/stevepur/transit-diffImage Difference images are used to confirm that an event is co-located with the target star, rather than being due to a nearby contaminant. Moving objects appear smeared out in difference images, while on-target microlensing events remain isolated.
@@
(3) RA/Dec event correlation: As a final check, we inspected the RA-Dec distribution of our candidates to search for extended trails as the objects move across the TESS field and cause brightening events on different stars. We identified multiple trails, all of which also corresponded to asteroids identified through our pixel-level analysis. Between our pixel-level and RA-Dec investigations, we rejected another 278 signals as likely asteroids.
@@
Note that we initially implemented the single-linkage clustering algorithm described by McDonald et al. (2021) to automatically flag these correlated structures. However, we found that the low relative stellar density of the TESS field of view with respect to K2 limited the efficacy of this method, and we removed it from our vetting procedure. In future work to further automate the process, we plan to instead incorporate shift-and-stack methods (e.g., Holman et al., 2019; Payne et al., 2019), which have already been used to detect asteroids with TESS FFI data (Woods et al., 2021).
@@ 3.5 Manual Inspection
We manually inspected both light curves and difference images for the surviving 122 candidates in order to confirm that the events appeared symmetric, unique, and consistent with our expectations for on-target lensing events. Most events that quickly failed visual inspection were obvious detrending artifacts caused by discontinuities in the QLP light curves, or were caused by stellar variability that fell beyond the thresholds of the cuts used during the automated vetting procedure. In some cases, inspection of the difference images revealed that a lensing candidate was offset from its true source. Such events were further investigated using the light curve of the correct source; all turned out to be variable stars.
@@
Based on flux- and pixel-level data from Sector 61 alone, we found two high-SNR signals associated with TIC-123147666 and TIC-107150013 that appeared to be possible microlensing signals. However, a single TESS sector is insufficient for identifying repeating events that occur on timescales near or longer than 27 days, such as periodic brightening events caused by tidal distortion in highly eccentric binary star systems (e.g., KOI-54; Welsh et al., 2011).
@@
We visually inspected all available TESS observations for our most promising candidates in search of repeated events, and found that TIC-123147666 featured similar brightening events in all other TESS sectors (Sectors 7, 8, 34, 35, and 62), consistent with a period of 26.2 days. This star was also previously identified as having a proper motion anomaly between the long-term proper motion vector and Gaia DR2 and HIPPARCOS measurements, indicative of the presence of a perturbing secondary object (Kervella et al., 2019). TIC-123147666 also has a high Renormalized Unit Weight Error (RUWE) value (8.10) according to Gaia DR3 observations (Gaia Collaboration et al., 2016, 2021), consistent with having a companion (Lindegren et al., 2018, 2021). We therefore classify TIC-123147666 as a false positive likely due to tidal interactions with a binary companion. Upon revisiting several of our other high-SNR, non-asteroid signals that had been manually rejected due to lack of consistency with lens models, we found that many were associated with repeating events throughout other TESS sectors. Future vetting efforts could be substantially improved and further automated by incorporating multi-sector data earlier in the process.
In contrast to TIC-123147666, we found no repeating events in other sectors for TIC-107150013 (Sectors 7 and 34), and the pre-detrended light curves had no indications of significant stellar variability or flares (Figure 2). Furthermore, its low RUWE value (0.99) is consistent with a single star, disfavoring the presence of an unresolved binary companion associated with this source.
It remains possible that a periodic signal may have fallen within the large data gaps between Sectors 7, 34, and 61; the planned 54 days of continuous observations of this star across TESS sectors 87 and 88 (2024 December 18 – 2025 February 11) will provide a future opportunity to rule out long-period signals. However, given the multiple lines of evidence that this star is not influenced by a binary companion, we take TIC-107150013 to be solitary and proceed with it as the source associated with our one remaining candidate lensing event.
Figure 2: Pre-detrended QLP light curves from Sectors 7, 34, and 61 for TIC-107150013, demonstrating that the microlensing candidate at 2987 days is a single, unique event based on the available TESS observations.
@@ 4 Results and Discussion
TIC-107150013 is located at a distance from Earth of $d_{s}=3193\pm 153$ pc based on parallax from Gaia DR3 (Table 1). Lacking a stellar radius from Gaia DR3, we adopt $R_{s}=12.91_{-1.51}^{+0.96}~{}R_{\odot}$ from Gaia DR2 (Gaia Collaboration et al., 2016, 2018). The star lies only $5.03^{\circ}$ below the Galactic equator, making it a relatively crowded region of the sky; however, at $T=13.10$ mag, TIC-107150013 is the brightest star within $1^{\prime}$. The microlensing candidate event occurs at $t_{\mathrm{BLS}}=2987.03$ days, was detected by BLS with a duration of $\tau=0.38$ days, and has a high SNR of 31 (Figure 3). In order to further assess the event’s candidacy, we analyzed the same signal using multiple light-curve extraction methods and sperformed dedicated lens modeling including second-order effects arising from the non-negligible angular extent of the source and limb-darkening.
@@
$\mu_{\alpha}$ $-0.395\pm 0.010$ Proper motion right ascension (mas yr-1) Gaia DR3
@@
Figure 3: Flux- and pixel-level data associated with the source star TIC-107150013, which features a high-SNR ($\mathrm{SNR}=31$) microlensing candidate event at 2987 days. The direct image reflects the typical field near the target (pink star), while the difference image reflects the pixels undergoing the largest brightening over the event. White circles show the locations of all stars down to 4 magnitudes fainter than the target star, with sizes scaled by brightness. The difference image is consistent with the lensing candidate being co-located with the target star.
4.1 Alternative Light Curve Extraction
QLP produces light curves using an aperture photometry approach combined with difference imaging. A reference image is constructed using the median combination of good quality images and this reference is subtracted from each frame to produce difference images. Aperture photometry is performed on the difference images to produce a difference flux time series, and the TESS magnitude provided in the TIC is used to scale the difference fluxes relative to the star’s average flux. The end result is a de-blended flux time series assuming any variations observed are associated with the target star (Huang et al., 2020). This approach relies on an accurate estimate of the star’s TESS magnitude, and the amplitude of the microlensing event will be inaccurate if the target star is not the true source.
To test the sensitivity of the interpretation of our candidate to an alternative method, we extracted a new Sector 61 light curve using the TESS-Gaia Light Curve (TGLC) procedure (Han & Brandt, 2023) with a $50\times 50$ pixel FFI cutout size. TGLC produces both aperture and point spread function (PSF) high-precision light curves corrected for contamination from nearby stars based on star positions and magnitudes from Gaia DR3. Here we adopt the TGLC aperture light curves, which offer more reliable amplitude estimates for dim stars in crowded fields like TIC-107150013 (Han & Brandt, 2023). The TGLC aperture light curves are produced by performing aperture photometry on a $3\times 3$ pixel aperture centred on the target star. To correct for contamination, TGLC estimates the median total flux of nearby stars in the Gaia DR3 catalog as well as the percentage of the target star’s light that falls on the aperture based on the shape of an effective PSF model. The aperture light curve’s median is shifted to the Gaia-predicted median multiplied by this percentage, resulting in a de-blended light curve corrected for the contamination of all known stars resolved by Gaia DR3.

### v2

@@ 3.3 Automated Flux-Level Vetting
A total of 305,446 signals met our basic detection criteria and were sent through a first round of automated vetting, which involved computing a set of metrics to be compared to pass-fail thresholds. If a signal failed to meet any of our criteria, it was rejected as a possible microlensing candidate. We do not claim that these empirically-determined thresholds are optimal and, as is the case for any vetting procedure, they may have removed true signals as well as false positives. However, they successfully produced a dataset of potential signals amenable to manual inspection (§3.5).
@@ 3.3.1 Initial Model Fitting
Each signal was first fit to three different models: a straight line parameterized by a slope and offset; a skew normal distribution parameterized by a central time, amplitude, scale ($\omega$), and shape ($\alpha$) (see left panel of Figure 2); and a simple point-source lens model parameterized by $t_{0}$, $u_{0}$, and $t_{E}$. Fitting was performed using the lmfit Python package (Newville et al., 2016). Data more than $2\tau$ from $t_{\mathrm{BLS}}$ were ignored in order to speed up the fitting process.
Figure 2: Examples of BLS-detected brightening events corresponding to flares (left) and scattered light (right). The flare event has a skew shape parameter $\alpha=137$ based on a fit to a skew normal distribution (red line), which indicates strong asymmetry. The scattered light event reflects the most common type of microlensing false positive in TESS data, representing $\sim$80% of BLS-detected events.
An event was rejected if a point lens was a worse fit compared to a straight line according to the Bayesian Information Criterion (BIC) of each fit, i.e., $\mathrm{BIC}_{\mathrm{lens}}>\mathrm{BIC}_{\mathrm{line}}$. A total of 9,993 signals failed this criteria.
@@ 3.3.2 Edge Effects and Gapped Events
BLS occasionally triggers at the edge of data gaps, especially if such edges correspond to rapid increases in brightness due to scattered light (see right panel of Figure 2), a well-known systematic effect within TESS data. [^1] https://archive.stsci.edu/missions/tess/doc/TESS_Instrument_Handbook_v0.1.pdf A number of events also lack a significant amount of data either during the brightening, or just before and/or after. To identify these cases, we computed the fraction of the light curve near the event that actually contained data. We found $n_{\mathrm{exp}}$, the number of data points expected within $2\tau$ of $t_{\mathrm{BLS}}$ given the 200-second cadence of the observations, as well as $n_{\mathrm{obs}}$, the number of data points actually observed over that time. We rejected 242,697 signals with $n_{\mathrm{obs}}/n_{\mathrm{exp}}<0.9$. Given that this cut rejected 79% of all vetted signals, edge effects and gapped events constituted the majority of TESS microlensing false positives.
@@ 3.3.3 Non-Unique Events
Microlensing events are intrinsically rare, arising from the chance alignment of source, lens, and observer. As such, the presence of multiple microlensing-like features at other times in a light curve suggest short-term periodic or quasi-periodic stellar variability, low-SNR noise/instrumental systematics, or some other false positive. This provides a means of mitigating such false positives by searching for unique features in a given light-curve. To quantify uniqueness, we adopted elements from the Model-Shift Uniqueness Test developed for Kepler transiting exoplanet candidate vetting (Coughlin, 2017). In summary, we computed an SNR time series by finding the SNR of events centered at each point in the light curve over the duration of the BLS-detected signal. We measured the significance of the primary signal ($\mathrm{SNR}_{\mathrm{pri}}$) as the largest SNR value in the SNR time series within a half event duration of the BLS detection. The most significant event at least two event durations from the primary was labeled the secondary event ($\mathrm{SNR}_{\mathrm{sec}}$), and the next most significant event at least two event durations from both the primary and secondary was labeled the tertiary event ($\mathrm{SNR}_{\mathrm{ter}}$). Finally, the most significant negative event ($\mathrm{SNR}_{\mathrm{neg}}$), representing the largest flux decrease at least three event durations from the primary and secondary, was also labeled.
We rejected signals for which the primary event was not sufficiently unique, as quantified by $\mathrm{SNR}_{\mathrm{pri}}<10$ (indicating that it is not unique compared to the noise) or $\Delta\mathrm{SNR}<2$ compared to the secondary, tertiary, or most negative events (indicating that it is not unique compared to other events at the same timescale). A total of 19,811 signals failed this criteria.
The ability of such a test to rule out false positives is limited due to TESS’s 27.4-day observation duration, as signals arising outside of this timespan are unobserved. This is in constrast to dedicated microlensing surveys, which rely on significantly longer baselines in order to establish the variable characteristics of source stars.
@@ 3.3.4 Flares
Flares are typically highly asymmetric brightening events characterized by a rapid increase in flux followed by a gradual decay, in contrast to the symmetric, bell-shaped peaks of microlensing events. We used the results of the skew normal distribution fit to identify this class of false positives. We rejected 205,230 signals with a skew shape of $|\alpha|>2$, suggesting strong asymmetry. Half of these also failed the edge effect/gapped event test, indicating that they were likely asymmetric events caused by scattered light. We also rejected 10,524 signals with a skew scale of $\omega<0.01$ days, as these are consistent with extremely short-duration flares.
@@ 3.4 Asteroid Rejection
A total of 31,885 signals passed all previous tests, and the most common remaining false positive scenario was brightening events caused by passing asteroids. We employed three major tests to reject asteroids: (1) cross-matching our candidates with known objects in the Solar System, (2) analyzing each event at the pixel level, and (3) searching for indications of moving bodies by matching correlated events across RA and Dec space.
@@
(1) Known asteroids: We queried the NASA/JPL Small-Body Identification API [^2] https://ssd-api.jpl.nasa.gov/doc/sb_ident.html for all known small bodies seen by TESS within the entire region in RA and Dec covered by Sectors 61–65 in each camera and CCD combination, extracting all objects brighter than 21 mag seen between 2962.79 and 3095.29 TESS BJD in steps of 2.5 days. We found 41,827 unique objects. Ephemerides for each were generated in 30-minute intervals using the JPL Horizons Solar System data and ephemeris computation service [^3] https://ssd.jpl.nasa.gov/horizons/. We then calculated the angular separation and difference in time between each asteroid and each surviving signal. We rejected 29,024 signals with angular separation of $0.07^{\circ}$ and time difference of <0.2<0.2 days.
@@
(2) Pixel-level analysis: Hundreds of asteroids were not rejected by this automated cut due to incompleteness of the Small-Body Database. We further removed asteroids by inspecting each event at the pixel level. Using $21\times 21$ pixel cutouts centered on each target star produced by tesscut (Brasseur et al., 2019), we computed the mean of all out-of-event images to represent the typical field near a source star when no brightening or dimming events are occurring on average. Out-of-event times were defined as those between $0.5\tau$ and $1.5\tau$ from the center of each event. We then subtracted the mean out-of-event image from each individual image in order to find the pixels undergoing the largest brightening at each point in time. We searched for changes in the locations of the brightest pixels within $1.5\tau$ of the event, which are consistent with an asteroid moving across the TESS pixels. Examples showing the maximum values of each pixel, for both an asteroid and a likely isolated signal, are provided in Figure 3, highlighting the clear difference between the two scenarios.
@@
Figure 3: Pixel-level data demonstrating typical signatures for asteroids (left) compared to isolated events (right), using examples associated with TIC-141479395 and TIC-432012378, respectively. These data were produced by subtracting the mean out-of-event image from each image within $1.5\tau$ of the center of a given event, and finding the maximum value of each pixel.
@@
We also produced difference images, employing the same techniques commonly used for transiting exoplanet vetting (Bryson et al., 2013), with the TESS difference imaging codebase transit-diffImage. [^4] https://github.com/stevepur/transit-diffImage Difference images are used to confirm that an event is co-located with the target star, rather than being due to a nearby contaminant. Nearby contaminants can include passing asteroids, other stars in neighboring pixels, or even stars in the same pixel as the target, given the large $21\times 21^{\prime\prime}$ size of TESS pixels. Moving objects will appear smeared out in difference images, while on- and off-target microlensing events will remain isolated.
@@
(3) RA/Dec event correlation: As a final check, we inspected the RA-Dec distribution of our events to search for extended trails as the objects move across the TESS field and cause brightening events on different stars. We identified multiple trails, all of which also corresponded to asteroids identified through our pixel-level analysis. Between our pixel-level and RA-Dec investigations, we rejected another 2861 signals as likely asteroids.
@@
Note that we initially implemented the single-linkage clustering algorithm described by McDonald et al. (2021) to automatically flag these correlated structures. However, we found that the low relative stellar density of the TESS field of view with respect to K2 Campaign 9 limited the efficacy of this method, and we removed it from our vetting procedure. In future work to further automate the process, we plan to instead incorporate shift-and-stack methods (e.g., Holman et al., 2019; Payne et al., 2019), which have already been used to detect asteroids with TESS FFI data (Woods et al., 2021).
@@ 3.5 Manual Inspection
We manually inspected both light curves and difference images for the surviving 452 candidates in order to confirm that the events appeared symmetric, unique, and consistent with our expectations for on-target lensing events. Most events that quickly failed visual inspection were obvious detrending artifacts caused by discontinuities in the QLP light curves, or were caused by stellar variability that fell beyond the thresholds of the cuts used during the automated vetting procedure. In some cases, inspection of the difference images revealed that the signal was offset from the target star. Such events were further investigated using the light curve of the correct source; all turned out to be variable stars.
@@
Based on flux- and pixel-level data, we found two high-SNR signals associated with TIC-123147666 and TIC-107150013 that had light-curve morphologies consistent with expectation of a short-duration microlensing event in Sector 61. However, a single TESS sector is insufficient for identifying repeating events that occur on timescales near or longer than 27 days, such as periodic brightening events caused by tidal distortion in highly eccentric binary star systems (e.g., KOI-54; Welsh et al., 2011).
@@
We visually inspected all available TESS observations for both of these signals in search of repeated events, and found that TIC-123147666 (Figure 4) featured similar brightening events in all other TESS sectors (Sectors 7, 8, 34, 35, 62, 87, and 88), consistent with a period of 26.2 days. This star was also previously identified as having a proper motion anomaly between the long-term proper motion vector and Gaia DR2 and HIPPARCOS measurements, indicative of the presence of a perturbing secondary object (Kervella et al., 2019). TIC-123147666 also has a high Renormalized Unit Weight Error (RUWE) value (8.10) according to Gaia DR3 observations (Gaia Collaboration et al., 2016, 2021), consistent with having a companion (Lindegren et al., 2018, 2021). We therefore classify TIC-123147666 as a false positive likely due to tidal interactions with a binary companion. Upon revisiting several of our other high-SNR, non-asteroid signals that had been manually rejected due to lack of consistency with lens models, we found that many were associated with repeating events throughout other TESS sectors. Future vetting efforts could be substantially improved and further automated by incorporating multi-sector data earlier in the process.
Figure 4: Detrended QLP light curves from Sector 61 (top) and Sector 62 (bottom) for TIC-123147666, which feature a brightening event that was initially identified in Sector 61 as a potential microlensing event before repeating events were identified in other TESS sectors. This event is likely caused by tidal distortion in a highly eccentric binary star system. The heartbeat nature of the event is more obvious in Sector 62, further emphasizing the need for multiple sectors when performing microlensing searches with TESS.
In contrast to TIC-123147666, we found no repeating events in other sectors for TIC-107150013 (Sectors 7, 34, 87, and 88). Furthermore, its low RUWE value (0.99) is consistent with a single star, disfavoring the presence of an unresolved binary companion associated with this source. There is potentially stellar variability at a timescale of order tens of days visible in the TESS data (Figure 5), though TESS is not well-suited to analysis of periodic variability longer than $\sim 13$ days due to systematics introduced by the telescope orbit (e.g. Hedges et al., 2020; Claytor et al., 2022). However, comparison to longer-baseline observations made by the long-running Optical Gravitational Lensing Experiment (OGLE) indicate that there is some degree of $\sim 40$-day periodic stellar variability. Mróz (2023) suggests that this may be due to the presence of stellar spots, indicating potential magnetic activity on this star. Additionally, Mróz (2023) presents two flares from the catalog compiled in Iwanek et al. (2019) that were qualitatively identified as being symmetric, and uses these events to make a rough estimate of the symmetric flare rate on giants in TESS data. However, as we show in §4, a Bayesian evaluation of which model is a better fit (between microlensing and flares) provides a quantitative means of more robustly addressing flare contamination in our data. We perform this analysis on the TIC-107150013 event as well as a dataset of flares with TESS’s fast cadence and high photometric precision to assess the level of contamination of symmetric flares in our microlensing search.
Figure 5: Pre-detrended QLP light curves from Sectors 7, 34, 61, 87, and 88 for TIC-107150013. There are no events with comparable significance to the event in Sector 61 at $\sim 2987$ days. The increase in noise in later sectors is due to their shorter cadences, which are associated with higher per-point uncertainties. Sector 7 was observed at 30-min cadence, Sector 34 was observed at 10-min, and Sectors 61–88 were observed at 200-sec. The light curves potentially indicate stellar variability over timescales on the order of tens of days, but TESS alone is not able to constrain the amplitude and period of this variability due to systematic differences between TESS orbits.
@@ 4 Results and Discussion
TIC-107150013 is located at a distance from Earth of $d_{s}=3193\pm 153$ pc based on parallax from Gaia DR3 (Table 1). Lacking a stellar radius from Gaia DR3, we adopt $R_{s}=12.91_{-1.51}^{+0.96}~R_{\odot}$ from Gaia DR2 (Gaia Collaboration et al., 2016, 2018a). The star lies $5.03^{\circ}$ below the Galactic plane, making it a relatively crowded region of the sky that is not very well explored by microlensing surveys; however, at $T=13.10$ mag, TIC-107150013 is the brightest star within $1^{\prime}$. The observed short-duration brightening event occurs at $t_{\mathrm{BLS}}=2987.03$ days, was detected by BLS with a duration of $\tau=0.38$ days, and has a high SNR of 31 (Figure 6). In order to further assess the event, we performed both microlensing fits and fits of stellar flare models to the data. We analyzed the same signal using multiple light-curve extraction methods and, in the case of the microlensing fit, performed dedicated lens modeling including second-order effects arising from the non-negligible angular extent of the source and limb-darkening.
@@
$\mu_{\alpha}$ $0.395\pm 0.010$ Proper motion right ascension (mas yr-1) Gaia DR3
@@
Figure 6: Flux- and pixel-level data associated with the source star TIC-107150013, which features a high-SNR ($\mathrm{SNR}=31$) transient signal at 2987 days. The direct image reflects the typical field near the target (pink star), while the difference image reflects the pixels undergoing the largest brightening over the event. White circles show the locations of all stars down to 4 magnitudes fainter than the target star, with sizes scaled by brightness. The difference image is consistent with the signal being co-located with the target star.
4.1 Alternative Light Curve Extractions
4.1.1 TESS-Gaia Light Curves
QLP produces light curves using an aperture photometry approach combined with difference imaging. A reference image is constructed using the median combination of good quality images and this reference is subtracted from each frame to produce difference images. Aperture photometry is performed on the difference images to produce a difference flux time series, and the TESS magnitude provided in the TIC is used to scale the difference fluxes relative to the star’s average flux. The end result is a de-blended flux time series assuming any variations observed are associated with the target star (Huang et al., 2020). This approach relies on an accurate estimate of the star’s TESS magnitude, and the amplitude of a microlensing event will be inaccurate if the target star is not the true source.
To test the sensitivity of the interpretation of the event to an alternative method of data reduction, we extracted a new Sector 61 light curve using the TESS-Gaia Light Curve (TGLC) procedure (Han & Brandt, 2023) with a $50\times 50$ pixel FFI cutout size. TGLC produces both aperture and point spread function (PSF) high-precision light curves corrected for contamination from nearby stars based on star positions and magnitudes from Gaia DR3. Here we adopt the TGLC aperture light curves, which offer more reliable amplitude estimates for dim stars in crowded fields like TIC-107150013 (Han & Brandt, 2023). The TGLC aperture light curves are produced by performing aperture photometry on a $3\times 3$ pixel aperture centred on the target star. To correct for contamination, TGLC estimates the median total flux of nearby stars in the Gaia DR3 catalog as well as the percentage of the target star’s light that falls on the aperture based on the shape of an effective PSF model. The aperture light curve’s median is shifted to the Gaia-predicted median multiplied by this percentage, resulting in a de-blended light curve corrected for the contamination of all known stars resolved by Gaia DR3.